OpenStudy (thefluffmuncher):

Need help finding the simplest radical form of 70

OpenStudy (misty1212):

do you know how to factor 70?

OpenStudy (thefluffmuncher):

no

OpenStudy (misty1212):

i don't believe you what number goes in to 70 evenly?

OpenStudy (thefluffmuncher):

10?

OpenStudy (misty1212):

ok so \[70=10\times 7\] is a start now can you factor \(10\)?

OpenStudy (thefluffmuncher):

5*2?

OpenStudy (misty1212):

right therefore \[70=2\times 5\times 7\] notice it has 3 factors, but none are squared

OpenStudy (thefluffmuncher):

ok

OpenStudy (misty1212):

so \[\sqrt{70}\] cannot be written any differently, it is already in simplest radical form

OpenStudy (thefluffmuncher):

oh

OpenStudy (thefluffmuncher):

thanks